# Arboretum: A Large Multimodal Dataset Enabling AI for Biodiversity

### v1

@@ 4 Models and Benchmarks
We now showcase demonstrate the utility of the Arboretum dataset by creating and benchmarking ArborCLIP, a new suite of vision-language foundation models for biodiversity.
4.1 Arboretum-40M
To start, we first construct Arboretum-40M, a subset comprising approximately 40M samples and 33K species, which we constructed by applying the data filtering tools described above.
The data included in this subset include all samples in the seven Arboretum categories posted on iNaturalist prior to January 27, 2024. We applied filtering criteria to exclude species with fewer than 30 images and imposed a maximum limit of 50,000 samples per species. We conducted semi-global shuffling and divided the data into mini-batches of approximately 50,000 samples each. From these mini-batches, 95% were randomly selected for training and validation, while the remaining 5% were reserved for testing. Detailed information can be found in Table 3.
Table 3: Training data sources used in Arboretum-40M and Diversity in Different Taxonomy Levels. We integrate taxonomic labels into the images.
Dataset Description Images Unique Classes TreeOfLife-10M Dataset combines a subset of iNaturalist, Encyclopedia of Life (EOL), BioScan-1M. 10.4M 454,103 Arboretum-40M One subset of Arboretum with size 40M. 39.9M 33,364 Level Uniques kingdom 3 phylum 14 class 50 order 311 family 1692 genus 11506 species 33364
@@ 4.2 New Benchmarks
From Arboretum, we create three new benchmark datasets for fine-grained image classification. In addition, we report results on several established benchmarks in the literature; see Table 4.
Arboretum-Balanced. To enforce a balanced species distribution across the seven categories, we curate Arboretum-Balanced. Each category includes up to 500 species, with 50 images per species. Table 6 in the Appendix shows each category’s exact number of species.
Arboretum-Unseen. To provide a robust benchmark for evaluating the generalization capability of models on unseen species, we curate Arboretum-Unseen. The test dataset was constructed by identifying species with fewer than 30 instances in Arboretum, ensuring that the dataset contains species that were unseen by ArborCLIP. Each species contained 10 images.
Arboretum-LifeStages. To assess the model’s ability to recognize species across various developmental stages, we curate Arboretum-LifeStages (see Figure 4). This dataset has 20 labels in total and focuses on insects, since these species often exhibits significant visual differences across their lifespan. Arboretum-LifeStages contains five insect species and utilized the observation export feature on the iNaturalist platform to collect data from 2/1/2024 to 5/20/2024 to ensure no overlap with the training dataset. For each species, life stage filters (egg, larva, pupa, or adult) were applied.
Figure 4: (a) Example images from Arboretum-Unseen. (b) Arboretum-Life-Stages with 20 class labels: four life stages (egg, larva, pupa, and adult) for five distinct insect species.
4.3 ArborCLIP: New vision-language foundation models for biodiversity
We use Arboretum-40M to train new CLIP-style foundation models, and then evaluate them on zero-shot image classification tasks. Following the implementation of Stevens et al. [35], we utilize a ViT-B/16 architecture initialized from the OpenAI CLIP weights [29], and train for 40 epochs. We also train a ViT-L/14 model from the MetaCLIP [41] checkpoint for 12 epochs, and a ViT-B/16 from the BioCLIP checkpoint for 8 epochs. All training hyperparameters are included in the Appendix (Section B). We compare with OpenAI’s ViT-B/16 CLIP model, the BioCLIP ViT-B/16 checkpoint, and MetaCLIP-CC ViT-L/14. We publicly release all code needed to reproduce our results here.
@@ 5 Experimental Results
Anim Birds 525 Scraped dataset of bird images from web search. [27] 89 8858988589\,88589 885 $525$ Taxonomic
@@
DeepWeeds Weed images collected in situ from eight rangelands across northern Australia [25] 17 5091750917\,50917 509 $9$ Common
@@
Metrics. For each of the benchmark datasets, we report top-1 zero-shot accuracy. For datasets where taxonomic information is available, we report accuracy over scientific names. For datasets where no taxonomic information is available, we utilize category names provided by the benchmark’s authors. For our Arboretum-Life-Stages benchmark, we consider a class label to be a unique (species | life-stage) tuple, resulting in a 20-class benchmark. Finally, we report an aggregate metric which is the weighted average over unique class labels across all benchmarks in our suite.
@@
Overview of results. In Table 5, we report the results of our core benchmark suite. At a high level, we observe that ArborCLIP variants achieve the best accuracy averaged over benchmarks. In particular, they perform extremely well on Arboretum-Balanced (a remarkable 91.1 top-1 accuracy over 2250+ class labels). ArborCLIP also does very well on the Fungi dataset (even though the Fungi class is not central to Arboretum-40M), and the DeepWeeds dataset. Therefore, ArborCLIP exhibits strong generalization capabilities across diverse datasets.
@@
We also observe that BioCLIP performs very well on Arboretum-Unseen and BioCLIP-Rare. The reasons might be that BioCLIP has seen approximately 450K species, and there might be nontrivial overlap with the species set in Arboretum-Unseen. On the other hand, it could be that ArborCLIP suffers from forgetting issues while training on Arboretum-40M. For BioCLIP-Rare, the dataset is a subset from EOL which BioCLIP did not see before, but TreeofLife contains the majority of the EOL dataset.
Limitations.
@@
We also evaluated all models on the challenging Confounding-species benchmark introduced in [4], but find that all models perform at or below random chance, and do not report results here; this could be an interesting avenue for follow-up work.
@@
In Table 7 in the Appendix, we report model performance at different levels of the taxonomic hierarchy. Generally, we find that models trained on web-scraped data perform better with common names, whereas models trained on specialist datasets perform better when using scientific names. Additionally, models trained on web-scraped data excel at classifying at the highest taxonomic level (kingdom), while models begin to benefit from specialist datasets like Arboretum-40M and Tree-of-Life-10M at the lower taxonomic levels (order and species). From a practical standpoint, this is not problematic: ArborCLIP is highly accurate at the species level, and higher-level taxa can be deterministically derived from lower ones.
Table 5: ArborCLIP performs well on a range of benchmarks. ArborCLIP-O was pretrained from an OpenAI model checkpoint, ArborCLIP-B from the BioCLIP checkpoint, and ArborCLIP-M from a MetaCLIP-cc checkpoint. AU stands for Arboretum-Unseen, using Scientific Names (n-classes=300). AB stands for Arboretum-Balanced, using Scientific Names (n-classes=2253). BCR stands for the BioCLIP Rare species benchmark (n-classes=400). F stands for Fungi (n-classes=25). I2 stands for Insects2 (n-classes=102). B stands for Birds525 (n-classes=525), LS for the Life Stages benchmark (n-classes=20), and DW for DeepWeeds (n-classes=9).
Dataset Arch AU AB BCR F I2 B LS DW Weighted Average
OpenAI ViT-B-16 12.9 7.3 10.9 11.5 10.2 50 56.5 10.3 14.7
BioCLIP ViT-B-16 68.2 62.2 30.2 45.1 20.8 68.7 18 19.9 58.5
MetaCLIP-cc ViT-L-14 24.9 15.4 20.5 24.6 16.1 70.1 64.3 14.7 25.0
ArborCLIP-O ViT-B-16 47.1 91.1 22.9 43.2 16.5 47.8 28 17 70.8
ArborCLIP-B ViT-B-16 53.8 82.2 23.7 53.9 16.9 57.1 15 18.4 67.2
ArborCLIP-M ViT-L-14 44.3 91.1 21.8 54.7 5.1 42.5 26.3 49.9 69.5
@@
Addressing these limitations will further enhance the applicability of models like ArborCLIP in real-world biodiversity monitoring tasks.
@@ 6 Concluding Discussion
We introduce Arboretum, the largest publicly accessible dataset designed to advance AI for biodiversity applications. This dataset, curated from the iNaturalist community science platform, includes 134.6 million images, surpassing existing datasets in scale by an order of magnitude. We anticipate that Arboretum will enable the development of AI models that can enable various digital tools ranging from pest control strategies, crop monitoring, and worldwide biodiversity assessment and environmental conservation.
We also believe that Arboretum can be used as a unique testbed for measuring progress on fine-grained image recognition. The success of ArborCLIP on Arboretum-Unseen underscores the importance of scaling up per-category sample size, or vertical scaling [8], in achieving high accuracy on long-tailed extreme-imbalance classification. However, BioCLIP continues to exhibit superior performance on several datasets, and we believe that this is because TreeofLife-10M contains an order-of-magnitude more classes (species) than Arboretum-40M. We invite the AI community to create new subsets of Arboretum with varying degrees of balance and species diversity, and use our tooling to measure model performance against current benchmarks.
@@ Acknowledgements
This work was supported by the AI Research Institutes program supported by the NSF and USDA-NIFA under AI Institute: for Resilient Agriculture, Award No. 2021-67021-35329. This was also partly supported by the NSF under CPS Frontier grant CNS-1954556. CH, BF, AN, and KM gratefully acknowledge the support of NYU IT High Performance Computing resources, services, and staff expertise.
@@ A.3 Scientific versus common names
Each image in the iNaturalist Open Dataset can be associated with its appropriate metadata through a group of four metadata CSV files,
representing photos, observations, taxa, and observers.
@@
The Arboretum metadata file used for model training contains the metadata columns phylum, class, order, family, genus, species, scientific_name, common_name for the seven Arboretum categories Aves, Arachnida, Insecta, Plantae, Fungi, Mollusca, and Reptilia. To ensure that only images and metadata from the seven Arboretum categories appear in our final dataset we use the taxa table to find the taxon in our categories then use it in a SQL query on the ancestry column of our aggregated metadata table.
The taxonomic rank columns are also found utilizing the ancestry metadata column. A difficulty in working with the ancestry metadata is present in that there is not a clear indication of what taxonomic rank a taxon id represents the ancestry string. This problem is exacerbated due to the presence of taxonomic ranks and dsub ranks whose presence is variable across different species. As such, a custom function is applied to each row to dynamically find the rank of each taxon id in the ancestry and then appropriately populate the taxon id to a metadata column of that rank. This process results in all taxonomies rank represented as metadata columns; only phyllum, class, order, family, genus and species are kept in the Arboretum metadata file.
The scientific name of a species is found using the name metadata column of our aggregated metadata table. The common name of a species is also useful metadata information. Unfortunately, the iNaturalist Open Data metadata files do not contain the common name information of a species. To address this, we curate a lookup table of the common names in our dataset. This is obtained from the iNaturalist Taxonomy DarwinCore Archive, Having obtained the common names for each species, we append it to the Arboretum-specific metadata.
A.6 Composition of Arboretum-40M
@@
Figure 5: Arboretum-40M Dataset Analysis: a) Consistent category distribution across Arboretum-40M and Arboretum-116M datasets. b) Species exhibit a long-tailed distribution. c) Impact of local vs. semi-global shuffling on species representation within training minibatches.
Table 6: Number of Unique Species in Each Category in Arboretum-Balanced.
@@
Appendix B ArborCLIP training details
We use Arboretum-40M to train new CLIP-style foundation models, and then evaluate them on zero-shot image classification tasks. Following the implementation of Stevens et al. [35], we utilize a ViT-B/16 architecture initialized from the OpenAI pretrained weights for our main model, and train for 40 epochs.

### v2

@@ 4 Models and Benchmarks
We now showcase and demonstrate the utility of the BioTrove dataset by creating and benchmarking BioTroveCLIP, a new suite of vision-language foundation models for biodiversity.
4.1 BioTrove-Train
BioTrove-Train is a curated subset comprising approximately 40M samples and 33K species, focused specifically on seven taxonomic categories: Aves, Arachnida, Insecta, Plantae, Fungi, Mollusca, and Reptilia. As discussed previously, the BioTrove dataset is accompanied by a flexible pipeline that enables users to apply customized filtering to select specific categories or subsets based on research needs, thereby allowing researchers to generate their own training datasets. For BioTrove-Train, these seven categories were pre-selected due to their significant impact on biodiversity and agricultural ecosystems, as well as their relative underrepresentation in standard image recognition models. Unlike megafauna, which are typically well-represented in existing models, these categories address unique challenges in biodiversity-focused AI.
This subset comprises data posted on iNaturalist prior to 2024-01-27. We applied strict filtering criteria to ensure high-quality data, excluding species with fewer than 30 images and capping the maximum number of images per species at 50,000. To maintain balance, we employed a semi-global shuffling method, organizing the data into mini-batches of approximately 50,000 samples. From these, 95% were randomly selected for training and validation, while the remaining 5% were reserved for testing. Detailed statistics can be found in Table 3.
Table 3: Training data sources used in BioTrove-Train and Diversity in Different Taxonomy Levels. We integrate taxonomic labels into the images.
Dataset Description Images Unique Classes TreeOfLife-10M Dataset combines a subset of iNaturalist, Encyclopedia of Life (EOL), BioScan-1M. 10.4M 454,103 BioTrove-Train One subset of BioTrove with size 40M. 39.9M 33,364 Level Uniques kingdom 3 phylum 14 class 50 order 311 family 1692 genus 11506 species 33364
@@ 4.2 New Benchmarks
We created three new benchmark datasets, all of which are non-overlapping curated subsets of the BioTrove dataset. These benchmarks focus on fine-grained image classification within the seven taxonomic categories: Aves, Arachnida, Insecta, Plantae, Fungi, Mollusca, and Reptilia. All benchmarks presented here are independent and strictly within these seven categories, without overlapping with each other or with the BioTrove-Train subset. Additionally, we report results on several established benchmarks from the literature (see Table 4).
BioTrove-Balanced. To ensure balanced species representation across the seven key taxonomic categories, we curate the BioTrove-Balanced benchmark. Each category includes up to 500 species, with 50 images per species, resulting in a total of 112,209 images. This balanced dataset provides a consistent foundation for model performance evaluations. The exact species counts for each category are detailed in Table 7 (see Appendix).
BioTrove-Unseen. To assess the ability of models to generalize to previously unseen species within the seven categories, we curated the BioTrove-Unseen benchmark. This dataset includes species from BioTrove-Train with fewer than 30 instances, ensuring they were unseen during training. Each species is represented by at least 10 images, with a total of 11,983 images. This benchmark tests the models’ robustness on rare species not encountered during training.
BioTrove-LifeStages. The BioTrove-LifeStages benchmark evaluates the model’s ability to recognize species across different developmental stages, focusing on insect species that exhibit significant visual variations throughout their life cycle. This dataset contains 20 labels representing four life stages (egg, larva, pupa, and adult) for five distinct insect species. The data was collected via the observation export feature on the iNaturalist platform between February 1, 2024, and May 20, 2024, ensuring no overlap with the training dataset. This benchmark allows for comprehensive evaluations of model performance across various life stages (see Figure 4).
Figure 4: (a) Example images from BioTrove-Unseen. (b) BioTrove-Life-Stages with 20 class labels: four life stages (egg, larva, pupa, and adult) for five distinct insect species.
4.3 BioTrove-CLIP: New vision-language foundation models for biodiversity
We use BioTrove-Train to train new CLIP-style foundation models and then evaluate them on zero-shot image classification tasks. Following the implementation of Stevens et al. [39], we utilize a ViT-B/16 architecture initialized from the OpenAI CLIP weights [33], and train for 40 epochs. We also train a ViT-L/14 model from the MetaCLIP [45] checkpoint for 12 epochs and a ViT-B/16 from the BioCLIP checkpoint for 8 epochs. All training hyperparameters are included in the Appendix (Section A.8). We compare with OpenAI’s ViT-B/16 CLIP model, the BioCLIP ViT-B/16 checkpoint, and MetaCLIP-CC ViT-L/14. We publicly release all code needed to reproduce our results here.
@@ 5 Experimental Results
Anim Birds 525 Scraped dataset of bird images from web search [31]. 89 8858988589\,88589 885 $525$ Taxonomic
@@
DeepWeeds Weed images collected in situ from eight rangelands across northern Australia [29]. 17 5091750917\,50917 509 $9$ Common
@@
Metrics. We evaluate model performance using top-1 zero-shot accuracy across all benchmark datasets. For datasets containing taxonomic information, we report accuracy based on scientific names, ensuring fine-grained classification. For datasets that lack explicit taxonomic details, we use the category labels as defined by the original benchmark authors. We compute an aggregate performance metric, which represents the weighted average accuracy over all unique class labels across the benchmark suite. This aggregate metric provides an overall view of model performance across diverse tasks.
To account for statistical variability, we include 95% confidence intervals for all reported metrics, calculated using the binomial proportion confidence interval method (denoted by ±). This provides a robust understanding of the performance and reliability of our results. As suggested during the review process, we incorporated this statistical analysis to strengthen the evaluation of our models.
@@
Overview of results. In Table 5, we report the results of our core benchmark suite. At a high level, we observe that BioTrove-CLIP variants achieve the best accuracy averaged over benchmarks. In particular, they perform extremely well on BioTrove-Balanced (a remarkable 91.1 top-1 accuracy over 2250+ class labels). BioTrove-CLIP also does very well on the Fungi dataset (even though the Fungi class is not central to BioTrove-Train), and the DeepWeeds dataset. Therefore, BioTrove-CLIP exhibits strong generalization capabilities across diverse datasets.
@@
We also observe that BioCLIP performs very well on BioTrove-Unseen and BioCLIP-Rare. The reasons might be that BioCLIP has seen approximately 450K species, and there might be nontrivial overlap with the species set in BioTrove-Unseen. On the other hand, it could be that BioTrove-CLIP suffers from forgetting issues while training on BioTrove-Train. For BioCLIP-Rare, the dataset is a subset from EOL which BioCLIP did not see before, but TreeofLife contains the majority of the EOL dataset.
@@
Limitations. We also evaluated all models on the challenging Confounding-species benchmark introduced in [4], but find that all models perform at or below random chance and do not report results here; this could be an interesting avenue for follow-up work.
@@
In Table 8 in the Appendix, we report model performance at different levels of the taxonomic hierarchy. Generally, we find that models trained on web-scraped data perform better with common names, whereas models trained on specialist datasets perform better when using scientific names. Additionally, models trained on web-scraped data excel at classifying at the highest taxonomic level (kingdom), while models begin to benefit from specialist datasets like BioTrove-Train and Tree-of-Life-10M at the lower taxonomic levels (order and species). From a practical standpoint, this is not problematic: BioTrove-CLIP is highly accurate at the species level, and higher-level taxa can be deterministically derived from lower ones.
Table 5: BioTrove-CLIP performance on various benchmarks. The top three rows are pre-trained checkpoints: OpenAI-B refers to OpenAI’s ViT-B-16 model, BioCLIP-B refers to the BioCLIP ViT-B-16 model, and MetaCLIP-L refers to the MetaCLIP-cc ViT-L-14 model. The bottom three rows are Biotrove-Clip models fine-tuned on different checkpoints: BT-Clip-O (from OpenAI-B), BT-Clip-B (from BioCLIP-B), and BT-Clip-M (from MetaCLIP-L). Benchmark abbreviations: BTU (Biotrove-Unseen, n=300), BTB (Biotrove-Balanced, n=2253), BCR (BioCLIP-Rare, n=400), F (Fungi, n=25), I2 (Insects-2, n=102), B (Birds-525, n=525), LS (Life-Stages, n=20), and DW (DeepWeeds, n=9). 95% confidence intervals (±) are included.
Model BTU BTB BCR F I2 B LS DW Weighted Avg.
OpenAI-B 12.9 ± 0.6 7.3 ± 0.15 10.9 ± 0.56 11.5 ± 1.98 10.2 ± 0.93 50.0 ± 0.33 56.5 ± 3.97 10.3 ± 0.45 14.7
BioCLIP-B 68.2 ± 0.83 62.2 ± 0.28 30.2 ± 0.82 45.1 ± 3.08 20.8 ± 1.25 68.7 ± 0.30 18.0 ± 3.07 19.9 ± 0.59 58.5
MetaCLIP-L 24.9 ± 0.77 15.4 ± 0.21 20.5 ± 0.72 24.6 ± 2.67 16.1 ± 1.13 70.1 ± 0.30 64.3 ± 3.83 14.7 ± 0.52 25.0
BT-CLIP-O 47.1 ± 0.89 91.1 ± 0.17 22.9 ± 0.75 43.2 ± 3.07 16.5 ± 1.14 47.8 ± 0.33 28.0 ± 3.59 17.0 ± 0.56 70.8
BT-CLIP-B 53.8 ± 0.89 82.2 ± 0.22 23.7 ± 0.76 53.9 ± 3.09 16.9 ± 1.15 57.1 ± 0.32 15.0 ± 2.86 18.4 ± 0.57 67.2
BT-CLIP-M 44.3 ± 0.89 91.1 ± 0.17 21.8 ± 0.74 54.7 ± 3.09 5.1 ± 0.68 42.5 ± 0.32 26.3 ± 3.52 49.9 ± 0.74 69.5
@@
Addressing these limitations will further enhance the applicability of models like BioTrove-CLIP in real-world biodiversity monitoring tasks.
@@ 6 Concluding Discussion
We introduce BioTrove, the largest publicly accessible dataset designed to advance AI for biodiversity applications. This dataset, curated from the iNaturalist community science platform, includes 161.9 million images, surpassing existing datasets in scale by an order of magnitude. We anticipate that BioTrove will enable the development of AI models that can enable various digital tools ranging from pest control strategies, crop monitoring, and worldwide biodiversity assessment and environmental conservation.
We also believe that BioTrove can be used as a unique testbed for measuring progress on fine-grained image recognition. The success of BioTrove-CLIP on BioTrove-Unseen underscores the importance of scaling up per-category sample size, or vertical scaling [9], in achieving high accuracy on long-tailed extreme-imbalance classification. However, BioCLIP continues to exhibit superior performance on several datasets, and we believe that this is because TreeofLife-10M contains an order-of-magnitude more classes (species) than BioTrove-Train. We invite the AI community to create new subsets of BioTrove with varying degrees of balance and species diversity and use our tooling to measure model performance against current benchmarks.
@@ Acknowledgements
We acknowledge support from the AI Research Institutes program supported by NSF and USDA-NIFA under AI Institute for Resilient Agriculture, Award No. 2021-67021-35329, and the NAIRR program for computing support.
@@ A.3 Scientific versus common names
Each image in the iNaturalist Open Dataset can be associated with its appropriate metadata through a group of four metadata CSV files, representing photos, observations, taxa, and observers.
@@
The BioTrove metadata file used for model training contains the metadata columns phylum, class, order, family, genus, species, scientific_name, common_name for the seven BioTrove categories Aves, Arachnida, Insecta, Plantae, Fungi, Mollusca, and Reptilia. To ensure that only images and metadata from the seven BioTrove categories appear in our final dataset we use the taxa table to find the taxon in our categories then use it in a SQL query on the ancestry column of our aggregated metadata table.
The taxonomic rank columns are also found utilizing the ancestry metadata column. A difficulty in working with the ancestry metadata is present in that there is not a clear indication of what taxonomic rank a taxon id represents the ancestry string. This problem is exacerbated due to the presence of taxonomic ranks and dsub ranks whose presence is variable across different species. As such, a custom function is applied to each row to dynamically find the rank of each taxon id in the ancestry and then appropriately populate the taxon id to a metadata column of that rank. This process results in all taxonomies rank represented as metadata columns; only phyllum, class, order, family, genus and species are kept in the BioTrove metadata file.
The scientific name of a species is found using the name metadata column of our aggregated metadata table. The common name of a species is also useful metadata information. Unfortunately, the iNaturalist Open Data metadata files do not contain the common name information of a species. To address this, we curate a lookup table of the common names in our dataset. This is obtained from the iNaturalist Taxonomy DarwinCore Archive, Having obtained the common names for each species, we append it to the BioTrove-specific metadata.
A.6 Composition of BioTrove and Related Datasets
In Table 6, we compare BioTrove with existing large-scale biodiversity datasets. BioTrove comprises 161.9 million research-grade images, representing approximately 372,966 species, and significantly surpasses other datasets in terms of both diversity and scale.
Table 6: Comparison of BioTrove with other biodiversity datasets.
Dataset BioTrove Wildlife Insights TreeOfLife BioScan iNaturalist 2017 [42] iNaturalist 2019 [16, 15] GBIF Backbone [8]
Size 161.9 million images 148.8 million images (52.6M wildlife images) 10.4M images 1.1M images 675,170 images 13.1M images 7.5M records
Diversity 366.6K species 3,682 species 454.1K species 8.3K species 5,089 species 166.8K species Millions of species
Labels Provided Common/scientific names, taxonomic hierarchies Species, location, timestamps, behavioral tags Common/scientific names, taxonomic hierarchies Scientific names, taxonomic ranks (family-species), DNA barcodes Common/scientific names, taxonomic ranks (genus-species) Common/scientific names, taxonomic ranks (genus-species) Species names, OTU identifiers
Data Source iNaturalist Open Dataset Camera traps, sensors iNaturalist, EOL, BioScan-1M Malaise trap specimens, DNA-barcodes iNaturalist iNaturalist Catalogue of Life, iBOL, UNITE, WoRMS, etc.
Key Features AI-ready pipeline, high-quality annotations, supports BioTrove-CLIP Automated processing, AI species recognition Rich hierarchical data, metadata, supports BioCLIP Insect-focused, high-resolution, taxonomic data, DNA codes Imbalanced classes, fine-grained taxonomy Large-scale species data, growth from 2017 Comprehensive taxonomy, cross-referencing datasets
AI-Ready Yes Yes Yes Yes No No No
A.7 Composition of BioTrove-Train
@@
Figure 5: BioTrove-Train Dataset Analysis: a) Consistent category distribution across BioTrove-Train and BioTrove-116M datasets. b) Species exhibit a long-tailed distribution. c) Impact of local vs. semi-global shuffling on species representation within training minibatches.
Table 7: Number of Unique Species in Each Category in BioTrove-Balanced.
@@
A.8 BioTrove-CLIP training details
We use BioTrove-Train to train new CLIP-style foundation models, and then evaluate them on zero-shot image classification tasks. Following the implementation of Stevens et al. [39], we utilize a ViT-B/16 architecture initialized from the OpenAI pretrained weights for our main model, and train for 40 epochs.